# XPrompt: Explaining Large Language Model’s Generation via Joint Prompt Attribution

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated impressive performances in complex text generation tasks. However, the contribution of the input prompt to the generated content still remains obscure to humans, underscoring the necessity of elucidating and explaining the causality between input and output pairs. Existing works for providing prompt-specific explanation often confine model output to be classification or next-word prediction. Few initial attempts aiming to explain the entire language generation often treat input prompt texts independently, ignoring their combinatorial effects on the follow-up generation. In this study, we introduce a counterfactual explanation framework based on joint prompt attribution, XPrompt, which aims to explain how a few prompt texts collaboratively influences the LLM’s complete generation. Particularly, we formulate the task of prompt attribution for generation interpretation as a combinatorial optimization problem, and introduce a probabilistic algorithm to search for the casual input combination in the discrete space. We define and utilize multiple metrics to evaluate the produced explanations, demonstrating both faithfulness and efficiency of our framework.
@@ 1 Introduction
Large Language Models (LLMs), such as GPT4 [1], LLaMA [2] and Claude [3], have shown excellent performance in various natural language generation tasks including question answering, document summarization, and many more. Despite the great success of LLMs, we still have very limited understanding on the LLM generation behavior – which parts in the input cause the model to generate a certain sequence. Unable to explain the causality between the input prompt and the output generation could cause failure in recognizing potential unintended consequences, such as harmful response [4, 5, 6, 7] and biased generation [8] attributed to a specific malicious description in the input. These issues undermine human trust in model usage, thus highlighting a pressing need for developing an interpretation tool that attributes how an input prompt leads to the generated content.
@@
Explaining LLM generation through prompt attribution involves extracting the most influential prompt texts on the model’s entire generation procedure, a realm that remains relatively under-explored in current research. While extensive works on input attribution are proposed for text classification interpretation [9, 10] and next-word generation rationale [11, 12], they can not be directly applied to explain the full generation sequence due to its complicated joint probability landscape and the autoregressive generation procedure. The complexity for interpreting LLM generation compounds as the model size increases.
Another line of works involves prompting LLMs to self-explain their behaviors [13]. This method relies on the model’s innate reasoning capabilities, although current findings suggest that these capabilities may not always be faithful [14, 15].
@@
There are limited existing attempts focusing on explaining the relationship between the input prompts and the complete generated sequence. The most relevant approach, Captum [16], sequentially determines the importance score for each token by calculating the variations in the joint probability of generating the targeted output sequence when the token is dropped from the model input. While being straightforard, this approach treats tokens as independent features, ignoring their joint semantic influence on the generated output. In fact, tokens may contain overlapping or complementary information. For example, given the input prompt: “Write a story about the doctor and his patient”, the most influential components are “doctor” and “patient”. Individually removing either of these words would not significantly alter the generated output, resulting in inaccurately low importance score for each of them. This is caused by the semantic interaction among these components, allowing the model to infer the meaning of the omitted one. Existing attribution methods that ablate each token individually fail to capture such combinatorial effect. A straightforward remedy might involve exhaustively assessing all possible combinations to observe the variations in the model generations, which however is impractical due to the vast search space with long-context input prompts.
To efficiently search the space for generating accurate prompt explanations, we develop our framework XPrompt, which provides the counterfactual explanation to highlight which components of input prompts have the fundamental effect on the generated context via solving a combinatorial optimization problem.
@@
We aim to explain the generation behavior of model outputs for any given prompt while take the joint effects of the prompt components into account. Assuming that removing the essential parts of the prompt would result in a significant variation in the model’s output, we propose the novel objective function and formulate our task of providing faithful counterfactual explanations for the input prompt as an optimization problem. To quantify the influence of token combinations in the prompt on the generations, we incorporate a mask approach for joint prompt attribution. Thus, our goal of extracting the explanations has been converted to finding the optimal mask of the input prompt. We solve this problem by a probabilistic search algorithm, equipped with gradient guidance and probabilistic updaten for efficient exploration in the discrete solution space. Our main contributions could be summarized as follows:
@@
  • We propose a general interpretation scheme for LLM generation task that attributes input prompts to the entire generation sequence. Notably, this recipe considers the joint influence of input token combinations on the generation. This motivation naturally formulates a combinatorial optimization problem for explaining generation with the most influential prompt texts.
@@
  • We demonstrate XPrompt, an efficient probabilistic search algorithm to solve the optimization problem. XPrompt works by searching better token combinations that lead to larger generation changes. It takes the advantage of both the gradient information and the probabilistic search-space strategy, thereby achieving an efficient prompt interpretation tool.
@@
  • Our framework demonstrates strong performance on language generation tasks including text summarization, question-answering, and general instruction datasets. The faithfulness of our explanations is evaluated based on a suit of comprehensive metrics considering generation probability, word frequency, and semantic similarity, verifying the transferability and effectiveness of our methods across a variety of tasks.
@@ 2 Related Work
While there are extensive works devoted to explaining language models in the context of text classification tasks [17, 18, 19], relatively few attempts [11, 12] are proposed to investigate the importance of prompt texts on the entire generation procedure especially for LLMs. This demonstrates a research gap that this work aims to fill in.
Explaining Language Generation
@@
There are many work explaining the predictions generated by LLMs by measuring the importance of the input features to the model’s prediction on the classification tasks. One group of studies perturb the specific input by removing, masking, or altering the input features, and evaluate the model prediction changes [20, 21]. The other group of works, such as integrated gradients (IG) [22] and first-derivative saliency [23], leverage the gradients of the output with respect to the input to determine the input feature importance.
@@
Moreover, a few studies utilize the surrogate model to explain the individual predictions of the black-box models, and the representative method is called LIME [24]. As for explaining the model’s generation behavior, Captum [16] calculates the token’s importance score by sequentially measuring the contribution of input token to the output, which lacks of the accounting for the semantic relationships between tokens. Another work, ReAGent [11], focuses on the next-word generation task, computing the importance distribution for the next token position. This method ignores the contextual dependencies in the the generated output and could not adequately account for dynamically changing generations. Our framework aims to interpret the joint effects of the input prompts on the entire output contexts with considering of the textual information covered the input prompt.
Self-explaining by Prompting
@@
As language models increase in scale, prompting-based models demonstrate remarkable abilities in reasoning [25], creativity [26], and adaptability across a range of tasks [27]. However, the complex reasoning processes of these models remain elusive and require tailored paradigm to better understand the prompting mechanism. For instance, the chain-of-thought (CoT) paradigm could explain the LLM behaviors by prompting the model to generate the reasoning chain along with the answers [13], as pre-trained LLMs have demonstrated a certain ability to self-explain their behaviors. However, recent studies have also suggested that the reasoning chain does not guarantee faithful explanations of the model’s behavior [28] and the final answer might not always follow the generated reasoning chain [14]. xLLM [29] enhances the fidelity of explanation derived from LLMs via a fidelity evaluator, which however is designed for classification tasks. Our efforts are concentrated on explaining LLM generation by analyzing the attribution of the input prompts to the output content, without depending on the model’s innate reasoning ability that are not yet satisfactory.
@@ 3 Preliminaries
We first introduce necessary notions depicting the LLM generation process, and then discuss limitations of prior attempts explaining the entire generation via prompt attribution.
Notions on LLM Generation
Denote a specific input prompt as a sequence of tokens $\bm{x}=(x_{1},\dots,x_{T})$, $x_{i}\in\{1,2,\dots,|V|\}$, where $|V|$ represents the vocabulary size, $T$ is the length of the input sequence and the set of all token indices is =$\mathcal{I}=\{1,2,\dots,T\}$. The corresponding generated output $\bm{y}$ could be represented as a sequence of tokens $\bm{y}=(y_{1},\dots,y_{S})$ with $y_{j}\in\{1,2,\dots,|V|\}$. The output tokens $\{y_{i}\}_{i=1}^{S}$ are generated from the LLM $f_{\bm{\theta}}$ parameterized by $\bm{\theta}$ in an autoregressive manner with the probability $p_{\bm{\theta}}(\bm{y}|\bm{x})$ as:
@@
There are limited prior works focus on explaining the relationship between the individual input prompts and the entire generated sequence, and their faithfulness is limited by treating input tokens independently. Specifically, Captum [16], calculates the importance score for each token by slightly perturbing the input $\bm{x}$ at the $i$-th token:
@@
These observations inspire us to develop a new prompt attribution method that considers the semantic correlation among tokens. Specifically, we assume that the content generated by LLMs is primarily influenced by a subset of tokens jointly. The remaining tokens serve as auxiliary or potentially less relevant information. Those subset of tokens, which fundamentally shape the model’s output, are viewed as the explanatory tokens for the generated content. In other words, explanatory tokens do not affect the model output independently, but jointly contribute to the generated responses.
@@ 4 Proposed Method
In this section, we propose XPrompt, a simple yet effective framework designed for generative tasks to explain the attribution of the input prompt by solving a discrete optimization problem. We start with formulating the general objective for the discrete optimization problem, and then we introduce our proposed probabilistic search algorithm for solving the problem.
Figure 1: Overview of XPrompt. Left: Demonstrating the pipline of the algorithm. Right: Illustrating the process of mask $\bm{m}$ sampling.
@@
In order to explain language generation via prompt attribution, instead of treating the prompt tokens independently as in previous works [16], we propose to evaluate the joint effect of $k$ prompt tokens on the generated sequence.
Specifically, consider a binary prompt mask $\bm{m}=(m_{1},\dots,m_{T})$ in which $m_{i}\in\{0,1\}$ indicates whether the $i$-th token is important or not. The joint probability of generating the original output sequence $\bm{y}$ given an input masked context $\bm{m}\odot\bm{x}$ is computed as $p_{\bm{\theta}}(\bm{y}|\bm{m}\odot\bm{x})$, where $\odot$ denotes the Hadamard product. In our paper, we aim to identify a binary mask $\bm{m}$ that maximizes the discrepancy in the probability of generating the same output $\bm{y}$ when comparing the masked input $\bm{m}\odot\bm{x}$ to the original input $\bm{x}$. A large discrepancy indicates that the masked token combination are the most influential components for generating $\bm{y}$, thus should jointly serve as the attributed explanation. Consequently, this involves training the binary mask $\bm{m}$ to optimize the following objective function:
$\max_{\bm{m}\in\{0,1\}^{T}}\mathcal{L}(\bm{m},\bm{x},\bm{y};\bm{\theta}):$ $=p_{\bm{\theta}}\left(\bm{y}|\bm{x}\right)-p_{\bm{\theta}}\left(\bm{y}|\bm{m}\odot\bm{x}\right)$ (3)
[…]
@@
Algorithm 1 Explainable Prompt Generator: XPrompt
@@
To tackle this challenge, we propose XPrompt, a novel explainable prompt generator, for efficiently obtaining an optimal solution for Eq. (3). We summarize our XPrompt in Algorithm 1. The high-level pipeline is illustrated in Figure 1 left: we initialize and maintain exactly $k$ entries in the mask $\bm{m}$ to be zero to enforce the constraint and capture their joint influence of being masked; the mask is then iteratively updated by sampling indexes for value swapping, searching towards the direction with increased generation loss $\mathcal{L}(\bm{m},\bm{x},\bm{y};\bm{\theta})$. At the core of this pipeline is the sampling and update of the discrete mask, which demands an efficient exploration in the vast search space. Figure 1 right shows this step, which is featured by the following two essential components, the gradient-guided masking and the probabilistic search update.
@@
Gradient-Guided Masking: Trivial solutions that set the mask $\bm{m}$ by uniformly random could cost massive sampling to hit the right optimization direction. Gradient is a common indicator of feature importance, as evidenced in existing practices [30, 6, 31]. Therefore, we propose to use gradient as a guidance to set and sample the mask for more efficient optimization. Specifically, we begin with the binary mask $\bm{m}^{(0)}=\bm{1}$ which indicates that all tokens in the input $\bm{x}$ are marked as non-explanatory ones. Then we compute the gradient $\nabla_{\bm{m}^{(0)}}\mathcal{L}(\bm{m}^{(0)},\bm{x},\bm{y};\bm{\theta})$ of the loss function in Eq. (3), and denote the magnitude of gradients as $\bm{g}=|\nabla_{\bm{m}^{(0)}}\mathcal{L}(\bm{m}^{(0)},\bm{x},\bm{y};\bm{\theta})|$. Note that the components with larger gradient magnitudes in $\bm{g}$ imply that altering the corresponding tokens could result in a sharper change to the generated outputs. In order to initialize the explanatory $k$ tokens guided by the gradient magnitude, we set the binary mask $\bm{m}^{(1)}$ where $m^{(1)}_{i}=0$ indicates $g_{i}$ is the top-$k$ value in $\bm{g}$. The gradient guidance and mask initialization are obtained following Line 1-2 in the algorithm.
Probabilistic Search Update: While gradient is informative, greedily determining explanatory tokens by top gradients lacks exploration, leading to suboptimal solutions. Therefore, we propose a probabilistic search mechanism for mask sampling and update. Specifically when updating the current binary mask $\bm{m}$, we iteratively sample a non-zero entry $l$ from the mask and swap its value with a sampled zero entry $v$ to explore a new (and potentially better) solution for the mask. Particularly, the sampling is also guided by the gradient calculated before: the non-zero entry in the mask (representing non-explanatory tokens) is sampled following probabilities calculated by the normalized gradient magnitudes, i.e., softmax($\bm{m}^{(n)}\odot\bm{g}$), and we employ a similar sampling strategy for zero entries (explanatory tokens). After swapping the mask indicators for the two sampled tokens $l$ and $v$, we generate a temporary mask $\bm{m}^{\text{tmp}}$. We then evaluate whether this temporary mask leads to a decrease in output probability. If it does, we update the binary mask to $\bm{m}^{\text{tmp}}$, otherwise we leave the binary mask as is. Consequently, without requiring intensive gradient computations, these sampling iterations keep discovering improved solutions for the discrete optimization problem shown in Eq. (3). We conclude this update by sampling process in Line 3-12.
While capturing the joint influence of being masked, the proposed XPrompt takes the advantage of both the gradient information and the search-space strategy, thereby achieving better efficiency than adopting either method alone. XPrompt requires only a single step of gradient calculation and avoids the need to convert between discrete and continuous masks. The obtained gradient information provides a favorable initial searching direction and reliable sampling probabilities that enhances the search efficiency.

### v2

# JoPA: Explaining Large Language Model’s Generation via Joint Prompt Attribution

@@ Abstract
Large Language Models (LLMs) have demonstrated impressive performances in complex text generation tasks. However, the contribution of the input prompt to the generated content still remains obscure to humans, underscoring the necessity of understanding the causality between input and output pairs. Existing works for providing prompt-specific explanation often confine model output to be classification or next-word prediction. Few initial attempts aiming to explain the entire language generation often treat input prompt texts independently, ignoring their combinatorial effects on the follow-up generation. In this study, we introduce a counterfactual explanation framework based on Joint Prompt Attribution, JoPA, which aims to explain how a few prompt texts collaboratively influences the LLM’s complete generation. Particularly, we formulate the task of prompt attribution for generation interpretation as a combinatorial optimization problem, and introduce a probabilistic algorithm to search for the casual input combination in the discrete space. We define and utilize multiple metrics to evaluate the produced explanations, demonstrating both the faithfulness and efficiency of our framework. Our code is available at  https://github.com/yuruic/JoPA.
@@ 1 Introduction
Large Language Models (LLMs), such as GPT4 (achiam2023gpt, ), LLaMA (touvron2023llama, ) and Claude (anthropic_2024_claude, ), have shown excellent performance in various natural language generation tasks including question answering, document summarization, and many more. Despite the great success of LLMs, we still have very limited understanding of the LLM generation behavior – which parts in the input cause the model to generate a certain sequence. Unable to explain the causality between the input prompt and the output generation could cause failure in recognizing potential unintended consequences, such as harmful response (dan, ; autodan-liu, ; gcg, ; autodan-zhu, ) and biased generation (wang2023decodingtrust, ) attributed to a specific malicious description in the input. These issues undermine human trust in model usage, thus highlighting a pressing need for developing an interpretation tool that attributes how an input prompt leads to the generated content.
@@
Explaining LLM generation through prompt attribution involves extracting the most influential prompt texts on the model’s entire generation procedure, a realm that remains relatively under-explored in current research. While extensive works on input attribution are proposed for text classification interpretation (chen2020learning, ; modarressi2023decompx, ) and next-word generation rationale (zhao2024reagent, ; vafa2021rationales, ), they can not be directly applied to explain the full generation sequence due to its complicated joint probability landscape and the autoregressive generation procedure. The complexity for interpreting LLM generation compounds as the model size increases. Another line of works involves prompting LLMs to self-explain their behaviors (wei2022chain, ). This method relies on the model’s innate reasoning capabilities, although current findings suggest that these capabilities may not always be faithful (Turpin2023LanguageMD, ; xu2024hallucination, ).
@@
There are limited existing attempts focusing on explaining the relationship between the input prompts and the complete generated sequence. The most relevant approach, Captum (captum, ), sequentially determines the importance score for each token by calculating the variations in the joint probability of generating the targeted output sequence when the token is dropped from the model input. While being straightforward, this approach treats tokens as independent features, ignoring their joint semantic influence on the generated output. In fact, tokens may contain overlapping or complementary information. For example, given the input prompt: “Write a story about the doctor and his patient”, the most influential components are “doctor” and “patient”. Individually removing either of these words would not significantly alter the generated output, resulting in inaccurately low importance score for each of them. This is caused by the semantic interaction among these components, allowing the model to infer the meaning of the omitted one. Existing attribution methods that ablate each token individually fail to capture such combinatorial effect. A straightforward remedy might involve exhaustively assessing all possible combinations to observe the variations in the model generations, which however is impractical due to the vast search space with long-context input prompts.
@@
To efficiently search the space for generating accurate prompt explanations, we develop our framework, Joint Prompt Attribution (JoPA), which provides a counterfactual explanation to highlight which components of input prompts have the fundamental effect on the generated context by solving a combinatorial optimization problem. We aim to explain the generation behavior of model outputs for any given prompt while taking the joint effects of the prompt components into account. Assuming that removing the essential parts of the prompt would result in a significant variation in the model’s output, we propose the novel objective function and formulate our task of providing faithful counterfactual explanations for the input prompt as an optimization problem. To quantify the influence of token combinations on the prompt on generations, we incorporate a mask approach for joint prompt attribution. Thus, our goal of extracting the explanations has been converted to finding the optimal mask of the input prompt. We solve this problem by a probabilistic search algorithm, equipped with gradient guidance and probabilistic updates for efficient exploration in the discrete solution space. Our main contributions can be summarized as follows:
@@
  • We propose a general interpretation scheme for LLM generation task that attributes input prompts to the entire generation sequence, considering the joint influence of input token combinations. This motivation naturally formulates a combinatorial optimization problem to explain generation with the most influential prompt texts.
@@
  • We demonstrate JoPA, an efficient probabilistic search algorithm to solve the optimization problem. JoPA works by searching better token combinations that lead to larger generation changes. It takes the advantage of both the gradient information and the probabilistic search-space strategy, thereby achieving an efficient prompt interpretation tool.
@@
  • Our framework demonstrates strong performance on language generation tasks including text summarization, question-answering, and general instruction datasets. The faithfulness of our explanations is assessed using metrics on generation probability, word frequency, and semantic similarity, confirming the methods’ transferability and effectiveness across tasks. Additionally, the explanations highlight the framework’s potential to enhance model safety and efficiency.
@@ 2 Related Work
While there are extensive works devoted to explaining language models in the context of text classification tasks (shi2022nearest, ; han2021ptr, ; shi2023chatgraph, ), relatively few attempts (zhao2024reagent, ; vafa2021rationales, ) are proposed to investigate the importance of prompt texts on the entire generation procedure especially for LLMs. This demonstrates a research gap that this work aims to fill in.
2.1 Explaining Language Generation
@@
There are many work explaining the predictions generated by LLMs by measuring the importance of the input features to the model’s prediction on the classification tasks. One group of studies perturb the specific input by removing, masking, or altering the input features, and evaluate the model prediction changes (10.1145/3461702.3462597, ; wu-etal-2020-perturbed, ). The other group of works, such as integrated gradients (IG) (sundararajan2017axiomatic, ), first-derivative saliency (li-etal-2016-visualizing, ), and mixed partial derivatives (tsang2020does, ) leverage the gradients of the output with respect to the input to determine the input feature importance. Although Archipelago (tsang2020does, ) explains the feature attributions by considering the combined effects of the input attributions, it targets the multilabel classification task and relies on the neutral baseline. As for the generation tasks, Diffmask learns the differentiable mask for each layer of the BERT model, but it aims to analyze how decisions are formed across network hidden layers by a simple probing classifier. In contrast, our framework targets on offering insights into the relationship between the input prompt and model generations by employing the mask to highlight the essential prompt attributions.
@@
Moreover, a few studies utilize the surrogate model to explain the individual predictions of the black-box models, and the representative method is called LIME (ribeiro2016should, ). As for explaining the model’s generation behavior, Captum (captum, ) calculates the token’s importance score by sequentially measuring the contribution of input token to the output, which lacks accounting for the semantic relationships between tokens. Another work, ReAGent (zhao2024reagent, ), focuses on the next-word generation task, computing the importance distribution for the next token position. This method ignores the contextual dependencies in the generated output and could not adequately account for dynamically changing generations. Our framework aims to interpret the joint effects of the input prompts on the entire output contexts with considering of the textual information covered the input prompt.
2.2 Self-explaining by Prompting
@@
As language models increase in scale, prompting-based models demonstrate remarkable abilities in reasoning (brown2020language, ), creativity (oppenlaender2022creativity, ), and adaptability across a range of tasks (khattak2023maple, ). However, the complex reasoning processes of these models remain elusive and require tailored paradigm to better understand the prompting mechanism. For instance, the chain-of-thought (CoT) paradigm could explain the LLM behaviors by prompting the model to generate the reasoning chain along with the answers (wei2022chain, ), as pre-trained LLMs have demonstrated a certain ability to self-explain their behaviors. However, recent studies have also suggested that the reasoning chain does not guarantee faithful explanations of the model’s behavior (jacovi-goldberg-2020-towards, ) and the final answer might not always follow the generated reasoning chain (Turpin2023LanguageMD, ). xLLM (chuang2024large, ) improves explanation fidelity from LLMs using a fidelity evaluator, tailored for classification tasks. We focus on explaining LLM generation by analyzing input prompt attributions to output content, independent of the model’s suboptimal reasoning abilities.
@@ 3 Preliminaries
We first introduce necessary notions for the LLM generation process, and then discuss limitations of prior attempts explaining the entire generation via prompt attribution.
LLM Generation Notions
Denote a specific input prompt as a sequence of tokens $\bm{x}=(x_{1},\dots,x_{T})$, $x_{i}\in\{1,2,\dots,|V|\}$, where $|V|$ represents the vocabulary size, $T$ is the length of the input sequence and the set of all token indices is $\mathcal{I}=\{1,2,\dots,T\}$. The corresponding generated output $\bm{y}$ could be represented as a sequence of tokens $\bm{y}=(y_{1},\dots,y_{S})$ with $y_{j}\in\{1,2,\dots,|V|\}$. The output tokens $\{y_{i}\}_{i=1}^{S}$ are generated from the LLM $f_{\bm{\theta}}$ parameterized by $\bm{\theta}$ in an autoregressive manner with the probability $p_{\bm{\theta}}(\bm{y}|\bm{x})$ as:
@@
There are limited prior works about explaining the relationship between the individual input prompts and the entire generated sequence, and their faithfulness is limited by treating input tokens independently. Specifically, Captum (captum, ), calculates the importance score for each token by slightly perturbing the input $\bm{x}$ at the $i$-th token:
@@
These observations motivate a new prompt attribution method that accounts for semantic correlations among tokens. We assume LLM outputs are primarily influenced by a subset of tokens acting jointly, while others serve as auxiliary information. These subset of tokens, which shape the output collectively, are identified as explanatory tokens, contributing to responses through joint effects rather than independently.
@@ 4 Proposed Method
In this section, we propose Joint Prompt Attribution (JoPA), a simple yet effective framework designed for generative tasks to explain the attribution of the input prompt by solving a discrete optimization problem. We start with formulating the general objective for the discrete optimization problem, and then we introduce our proposed probabilistic search algorithm for solving the problem.
Figure 1: Overview of JoPA. Left: Demonstrating the pipeline of the algorithm. Right: Illustrating the process of mask $\bm{m}$ sampling.
@@
In order to explain language generation via prompt attribution, instead of treating the prompt tokens independently as in previous works (captum, ), we propose to evaluate the joint effect of $k$ prompt tokens on the generated sequence.
Specifically, consider a binary prompt mask $\bm{m}=(m_{1},\dots,m_{T})$ in which $m_{i}\in\{0,1\}$ indicates whether the $i$-th token is important or not. The joint probability of generating the original output sequence $\bm{y}$ given an input masked context $\bm{m}\odot\bm{x}$ is computed as $p_{\bm{\theta}}(\bm{y}|\bm{m}\odot\bm{x})$, where $\odot$ denotes the Hadamard product. In our paper, we aim to identify a binary mask $\bm{m}$, which is a (discrete) learnable parameter, that maximizes the discrepancy in the probability of generating the same output $\bm{y}$ when comparing the masked input $\bm{m}\odot\bm{x}$ to the original input $\bm{x}$. A large discrepancy indicates that the masked token combination are the most influential components for generating $\bm{y}$, thus should jointly serve as the attributed explanation. Consequently, this involves training the binary mask $\bm{m}$ to optimize the following objective function:
$\max_{\bm{m}\in\{0,1\}^{T}}$ $\mathcal{L}(\bm{m},\bm{x},\bm{y};\bm{\theta})$ (3)
$=p_{\bm{\theta}}\left(\bm{y}|\bm{x}\right)-p_{\bm{\theta}}\left(\bm{y}|\bm{m}\odot\bm{x}\right)$
$\text{s.t.}\quad|\bm{m}|_{1}$ $=T-k,$
@@
Algorithm 1 Explainable Prompt Generator: JoPA
@@
2:Set $\bm{m}^{(1)}$ as the $\text{top-}k$ value mask for $\bm{g}$: $\bm{m}^{(1)}=\bm{m}^{(0)};m_{i}^{(1)}=0,\forall i\in\text{argtop-}k(\bm{g})$
@@
To tackle this challenge, we propose JoPA, a novel explainable prompt generator, for efficiently obtaining an optimal solution for Eq. (3). We summarize our JoPA in Algorithm 1. The high-level pipeline is illustrated in Figure 1 left: we initialize and maintain exactly $k$ entries in the mask $\bm{m}$ to be zero to enforce the constraint and capture their joint influence of being masked; the mask is then iteratively updated by sampling indexes for value swapping, searching towards the direction with increased generation loss $\mathcal{L}(\bm{m},\bm{x},\bm{y};\bm{\theta})$. At the core of this pipeline is the sampling and update of the discrete mask, which demands an efficient exploration in the vast search space. Figure 1 right shows this step, which is featured by the following two essential components, the gradient-guided masking and the probabilistic search update.
@@
Gradient-Guided Masking: Trivial solutions that set the mask $\bm{m}$ by uniformly random could cost massive sampling to hit the right optimization direction. Gradient is a common indicator of feature importance, as evidenced in existing practices (ebrahimi2018hotflip, ; gcg, ; shin-etal-2020-autoprompt, ). Therefore, we propose to use gradient as a guidance to set and sample the mask for more efficient optimization. Specifically, we begin with the binary mask $\bm{m}^{(0)}=\bm{1}$ which indicates that all tokens in the input $\bm{x}$ are marked as non-explanatory ones. Then we compute the gradient $\nabla_{\bm{m}^{(0)}}\mathcal{L}(\bm{m}^{(0)},\bm{x},\bm{y};\bm{\theta})$ of the loss function in Eq. (3), and denote the magnitude of gradients as $\bm{g}=|\nabla_{\bm{m}^{(0)}}\mathcal{L}(\bm{m}^{(0)},\bm{x},\bm{y};\bm{\theta})|$. Note that the components with larger gradient magnitudes in $\bm{g}$ imply that altering the corresponding tokens could result in a sharper change to the generated outputs. In order to initialize the explanatory $k$ tokens guided by the gradient magnitude, we set the binary mask $\bm{m}^{(1)}$ where $m^{(1)}_{i}=0$ indicates $g_{i}$ is the top-$k$ value in $\bm{g}$. The gradient guidance and mask initialization are obtained following Line 1-2 in the algorithm.
[…]
JoPA combines gradient information with a search-space strategy, achieving greater efficiency than using either alone. It requires only one gradient calculation step and avoids converting between discrete and continuous masks. Gradients offer an efficient search direction and reliable sampling probabilities. Theoretical insights are detailed in Appendix A.1.